What is the electric potential energy of the elect
Answer #1

given

d = 0.4*10^(-9) m

L = 1.5*10^(-9) m

the potential energy of the electron

U = V1 * q + V2 * q

U = kQcos\Theta/L + kQcos\Theta /L

U = 2(kQcos\Theta /L )   

where cos\Theta  = 0.96 { as tan\Theta = d/L}

U = 2(9*109 * 1.6*10-19* 0.96 * 1.6*10-19 /1.5*10-9)

U = 1.536 *10-19 J

answer